OpenStudy (anonymous):

1. P (-1.98 is less than or equal to z less than or equal to 0.49) Please help me on how to split this and solve...it's related to normal random variable.

ganeshie8 (ganeshie8):

\(\large P(-1.98 \le z \le 0.49) = P(z=0.49) - P(z=-1.98)\)

ganeshie8 (ganeshie8):

know how to interpret zscore tables ?

ganeshie8 (ganeshie8):

http://www.math.upenn.edu/~chhays/zscoretable.pdf

ganeshie8 (ganeshie8):

\(\large P(z=0.49) = 0.6879\) \(\large P(z=-1.98) = 0.0239\)

ganeshie8 (ganeshie8):

subtract them both

OpenStudy (anonymous):

thank you ganeshie8!!!

ganeshie8 (ganeshie8):

u wlc :)